Data concerning Mika Lumber Company's single product appear below:

elling price per unit$ 150.00

Variable expense per unit$ 34.50Fixed

expense per month$ 466,620

The break-even in monthly unit sales is closest to: (Round your intermediate calculations to 2 decimal places.)

6,892

3,111

4,040

13,525
Answer #1

Contribution margin=Sales-Variable cost  

=(150-34.5)=$115.5 per unit

Breakeven=Fixed expenses/Contribution margin

=466,620/115.5

=4040 units